QUIZ NAVIGATION Question 8 Two train cars filled with soy beans escape from the Hertage Cooperative in Ada and roll along the (very flat and straigh) rairoad track east towards Dola. The frst has a mass of 2.00 x 10 kg and moves at 3.00 m/s to the right. The second has a mass of 1.00 x 10 kg and moves at 6.00 ms to the right. When the second car catches the first, the two cars collide and couple together (ie. the collision is perfectly inelastic). what is the velocity after the collision? hint Assume that the net force on the system of the two cars is zero during the collision and therefore you can apply the conservation of momentum for the system of the two cars. 1 2 3 4 5 Points out of 100 car r Remove fiag Finish attempt Time left 0:03:30 Select one: O a. 4.00 m/s to the right o b. 3.00 m/s to the left c. 4.00 m/s to the left d. Zero e.3.00 m/s to the right
media%2F2c4%2F2c4f014b-4797-4e9b-90ec-fc
media%2F189%2F18956e60-997c-4a63-a023-96
Answer #1

Question 8:

Using conservation of momentum.

Initial momentum of the system is given by

P_{initial} = m_1v_1+m_2v_2 = 2*10^4*3 + 1*10^4*6 = 12*10^4 kgm/s

When trains collide and get attached, let the velocity is v, then final momentum is given by

P_{final} = (m_1+m_2)v = 3*10^4*v

Now, initial and final momentum should be equal. Thus,

3*10^4*v = 12*10^4kgm/s \\ \Rightarrow v = 4m/s

SO, both trains move with 4m/s towards right. Answer is option a.

Question 7:

acceleration is along the ground. Thus, component of force along the ground is given by

F_g = F*cos(60) = 112*\frac{1}{2} = 56N \\ \\ F_g = ma \\ \Rightarrow a = \frac{F_g}{m} = \frac{56}{25} = 2.25 m/s^2

Normal force exerted by ground is given by

N = mg-Fsin(60) = 25*9.8-112*\frac{\sqrt{3}}{2} = 245-97 \\N= 148N

Normal force acts upwards.

SO answer is option d.

Question 6:

Since the desk did not move, acceleration of the desk is zero.

Frictional force counteracts your force, thus frictional force = 99.5N towards left.

So, option b is correct